OpenStudy (anonymous):

Help me find the equation of the line passing through these points. (2,17), (5,41)

OpenStudy (anonymous):

to get you started, first you want to find the slope of the line and then you can plug that into the formula to find the equation of a line

OpenStudy (anonymous):

do you remember the formula to find slope?

OpenStudy (anonymous):

y^2-y^1 over x^2-x^1

OpenStudy (anonymous):

y1-y2/x1-x2 = slope

OpenStudy (anonymous):

yes, that is it

OpenStudy (anonymous):

what is the slope?

OpenStudy (anonymous):

24?

OpenStudy (anonymous):

I get 8 for the slope

OpenStudy (anonymous):

you should get 24 on top and 3 on the bottom which is equal to 8

OpenStudy (anonymous):

wow...... i understand now i was saying 5-2 was 2 when really it 3

OpenStudy (anonymous):

alright so now that you have the slope, you want to use the formula y-y1=m(x-x1) to find the equation of a line

OpenStudy (anonymous):

m=slope

OpenStudy (anonymous):

so the answer would be y=8x+1?

OpenStudy (anonymous):

because then you must multiply that 8 by either one of the pairs and i picked 2,17 so 8 x 2+ 16 and minus 17 equals 1

OpenStudy (anonymous):

I got y = 8x - 11